OpenStudy (anonymous):

Rectangle EFGH's base is b, its height is h, and it has a perimeter of 48 units. Determine the area of the rectangle for b=6, b=8, b=12, and b=16. Based on these results, what conjecture can you make about maximizing the area of a rectangle with a fixed perimeter?

OpenStudy (lgbasallote):

go @jhonyy!! hehehe

jhonyy9 (jhonyy9):

perimeter 48 so when b=6 than the height will be (48-12)/2 =36/2 =18 so than the area will be 18*6=108 - when b=8 then the height will be 48-16=32/2=16 so than the area will be 16*6=96 - when b=12 --- h=(48-24)/2=24/2=12 --- area=12*6=72

jhonyy9 (jhonyy9):

easy boy !!!

jhonyy9 (jhonyy9):

so with out he-he !!!

OpenStudy (anonymous):

Whoa, thanks!

jhonyy9 (jhonyy9):

stefaniee ok ?

jhonyy9 (jhonyy9):

good luck bye